you need to express the problem as an exponential function whose value is the percentage of the remaining thorium after x million years, and then use the function and the following two properties to find the age of meteorite in million years:
The level of thorium in a sample decreases by a factor of one-half every 2 million years. A meteorite is discovered to have only 8.6% of its original thorium remaining. How old is the meteorite? Thank you

Answer by stanbon(75887)

The level of thorium in a sample decreases by a factor of one-half every 2 million years. A meteorite is discovered to have only 8.6% of its original thorium remaining. How old is the meteorite?
-----------
A(t) = Ao*(1/2)^(t/2*10^6)
-----------------------
0.086Ao = Ao*(1/2)^(t/(2 million))
(1/2)^(t/2 million) = 0.086
Take the log to get:
(t/2 million) = log(0.086)/log(1/2) = 3.53951953
t = (2 million)*(3.53951953)
t = 7,079,039.06
=====================
Cheers,
Stan H.